# Efficient Individually Rational Recommender System under Stochastic Order

### v1

@@ 2.2 Stochastic Order
Assumption 2.1.
The reward of the arms in $\textnormal{neg}(A)$ are stochastically ordered.
Notice that this assumption applies only to a priori inferior arms, i.e., the arms in $\textnormal{neg}(A)$, and not to all arms. There are many natural cases for this assumption: Unit-variance (or any fixed variance) Gaussian as in Example 2.1, Bernoulli, log-normal, truncated normal, etc. When a formal statement relies on Assumption 2.1, we mention it explicitly.
@@ 3 An Auxiliary Goal Markov Decision Process
3.1 Motivation for GMPD: Bernoulli Trials
@@
We call mixing an observed arm with a positive realized reward and an unobserved arm with a negative expected reward a Bernoulli Trial. Observation 3.1 suggests that if we discover an arm with a positive reward, we can use Bernoulli trials to reveal all arms in finite time almost surely.
To illustrate, recall Example 2.1. Assume that we are in the third round, have already explored $a_{1}$ and $a_{2}$, and discovered that $X(a_{1})=x(a_{1})>0$ and that $X(a_{2})<0$. Furthermore, we have not yet explored $a_{3}$ and $a_{4}$ (one of which may still hide the highest realized reward among all arms). Consider the portfolio
@@
thus, $\boldsymbol{p}$ is IR. If we pick this portfolio, Nature flips coins to choose either $a_{1}$ and $a_{3}$, with the latter being selected with positive probability. This Bernoulli trial might result in selecting $a_{1}$, but we can repeat it until Nature picks $a_{3}$. The number of rounds required to explore $a_{3}$ follows the Geometric distribution with a success probability of $\nicefrac{{x(a_{1})}}{{x(a_{1})-\mu(a_{3})}}$ in each Bernoulli trial. Indeed, by executing Bernoulli trials until the first success, we can guarantee exploring $a_{3}$. Zooming out from Example 2.1 to the general case, Observation 3.1 suggests that once a positive reward is realized, we can execute enough Bernoulli trials to explore all the arms in finite time.
Observation 3.1 calls for modeling that abstracts the setting once a positive reward is realized. More specifically, we can focus on the case where all unexplored arms are revealed instantly after a positive reward is discovered. To that end, we introduced the induced constrained Goal Markov Decision Process (GMDP), which we present in the next subsection. Our goal is to find the optimal policy for this GMDP, and then translate it to an asymptotically optimal algorithm for the corresponding IRSR instance. [^4] We keep the terms mechanism and algorithm for solutions to IRSR and use the term policy for solutions of the GMDP.
3.2 An Auxiliary Goal Markov Decision Process
@@
  • In every state $s$, the feasible actions are the set of IR portfolio w.r.t. the prior information, i.e., from $\textit{IR}(s)=\left\{\boldsymbol{p}\in\Delta(s):\sum_{a\in s}\boldsymbol{p}(a)\mu(a)\geq 0\right\},$ where $\Delta(s)$ is the set of all distributions over the elements of $s$. If $\textit{IR}(s)$ is empty, then we say that $s$ is terminal.
@@
    $R(s)\stackrel{{\scriptstyle\text{def}}}{{=}}\begin{cases}\max_{a\in A}X(a)&\textnormal{if $\textit{IR}(s)=\emptyset$ and $\max_{a^{\prime}\in A\setminus s}X(a^{\prime})>0$}\\ 0&\textnormal{otherwise}\end{cases}.$
@@
A (stationary) policy is a mapping from states to portfolios. Due to standard arguments, which are further elaborated in Section A, there exists an optimal policy that is also stationary. Hence, we limit our attention to stationary policies only. Let $W(\pi,s)$ be the state-value function: The expected reward of a policy $\pi$ starting at the initial state $s$. Namely,
$W(\pi,s)=\begin{cases}R(s)&\textnormal{if }\textit{IR}(s)=\emptyset\\ \sum_{a\in s}\Pr(s\setminus\{a\}\mid s,\pi(s))W(\pi,s\setminus\{a\})&\textnormal{otherwise}\end{cases}.$ (3)
@@ 3.3 $\mathcal{P}$-valid Policies and an Inefficient Solution via Dynamic Programming
mixes $a_{i}$ and $a_{j}$ while maximizing the probability of exploring $a_{j}$ (the a priori inferior arm). The reader can verify that $\boldsymbol{p}_{i,j}$ is indeed IR, yielding an expected value of precisely zero. For completeness, for every $a_{i}\in\textnormal{pos}(A)$, we also define $\boldsymbol{p}_{i,i}$ as a deterministic selection of $a_{i}$, e.g., $\boldsymbol{p}_{i,i}(a)=1$ if $a=a_{i}$, and zero otherwise. Next, we define $\mathcal{P},\mathcal{P}^{\prime}$ such that
@@
Definition 3.1 ($\mathcal{P}$-valid Portfolio).
A portfolio $\boldsymbol{p}$ is $\mathcal{P}$-valid with respect to a state $s\in\mathcal{S}$, if
  • $\boldsymbol{p}\in\textit{IR}(s)$, and
@@
We say a policy is $\mathcal{P}$-valid if, for all states $s\in\mathcal{S}$, $\pi(s)$ is a $\mathcal{P}$-valid portfolio. Notice that $\mathcal{P}$-valid policies are IR by definition. Furthermore, they mix at most two actions from $A$ in each state; hence, there are $O(K^{2})$ actions $\mathcal{P}$-valid policies can take in every state. This is in sharp contrast to the set of actions $\textit{IR}(s)$ for a state $s$, which is generally a convex polytope with infinitely many actions.
Due to the linearity of the value function $W(\pi,s)$ in $\pi(s)$, the GMDP exhibits a nice structural property, as captured by the following Proposition 3.1.
Proposition 3.1.
@@
The proof of this proposition appears in Section H. While the GMDP is defined using a continuous action space in each state (the convex polytope $\textit{IR}(s)$), Proposition 3.1 allows us to narrow down the search to $\mathcal{P}$-valid portfolios. Particularly, we can find an optimal policy with dynamic programming using a bottom-up approach in time $O(2^{K}K^{2})$, as we explain next.
@@ 3.4 Efficiently Computing an Optimal GMDP Policy
Input: a state $s\subseteq A$
Output: a $\mathcal{P}$-valid portfolio or $\emptyset$
1 if $s$ is terminal then
2       return $\emptyset$
3else
4       pick any arbitrary $a_{i}\in\textnormal{pos}(s)$
5       if $\textnormal{neg}(s)=\emptyset$ then
6             return $\boldsymbol{p}_{i,i}$
7
8      else
9             pick $a_{j^{\star}}\in\operatorname*{arg\,max}_{a_{j}\in\textnormal{neg}(s)}\mu(a_{j})$
10             return $\boldsymbol{p}_{i,j^{\star}}$
11
@@
We start this subsection by presenting an optimal GMDP policy (hereinafter OGP for shorthand), which we formalize via Policy 1, and argue for its optimality later. Given a state $s$, ${\small\textnormal{{OGP}}}(s)$ operates as follows. If $s$ is terminal, it returns the empty set (Lines 11). Otherwise, if $s$ is non-terminal, we enter the “else” clause in Line 1, and pick any arbitrary arm $a_{i}$ from $\textnormal{pos}(s)$. Then, we have two cases. If $\textnormal{neg}(s)$ is empty, OGP returns $\boldsymbol{p}_{i,i}$ (Line 1). Else, it picks the best arm from $\textnormal{neg}(s)$ in terms of expected reward, which we denote by $a_{j^{\star}}$ (Line 1), and returns $\boldsymbol{p}_{i,j^{\star}}$ (Line 1). Overall, OGP returns $\emptyset$ if $s$ is terminal or a $\mathcal{P}$-valid portfolio if $s$ is non-terminal.
Theorem 3.1.
Fix any arbitrary IRSR instance satisfying Assumption 2.1 and observe its induced GMDP instance. For any state $s\subseteq A$, OGP prescribed in Policy 1 satisfies $W({\small\textnormal{{OGP}}},A)=W^{\star}(A)$.
  • Proof Overview of Theorem 3.1.We defer the full proof to Section A and outline it below. To prove this theorem, we focus on the dynamic programming explained at the end of the previous subsection. We unveil its crux and flesh out the index structure OGP uses. A crucial ingredient of the analysis is the probability of reaching the empty state, representing the case in which we explored all arms. More formally, for every state $s\in\mathcal{S}$ and a policy $\pi$, we denote by $Q(\pi,s)$ the probability starting at $s\subseteq A$, following the policy $\pi$ and reaching the empty state. Note that, similarly to $W$ the state-value function, $Q$ is defined recursively: Namely, if $\pi(s)=\boldsymbol{p}_{i,j}$ for a non-terminal state $s$, then
@@
    Lemma 3.1 (Equivalence Lemma).
@@
    The Equivalence Lemma asserts that $Q$ is policy-independent, a property that significantly simplifies the analysis. The proof of this lemma is intricate and relies on a careful factorization of the elements of $Q$. Interestingly, the arguments we have mentioned so far, including the Equivalence Lemma, are independent of Assumption 2.1.
    Equipped with the Equivalence Lemma, we use the recursive structure of $Q$ and the reward function $W$ to prove the optimality of OGP inductively over the sizes of $\textnormal{pos}(s)$ and $\textnormal{neg}(s)$. This step makes use of further insights into the dynamic programming procedure, such as monotonicity in the set of available arms. ∎
@@
Figure 1: Illustration of OGP for Example 2.1.
@@ 3.5 Stochastic Order and Non-triviality
Before leveraging OGP to an optimal IRSR algorithm, it is important to describe why Theorem 3.1 is non-trivial to prove, even with the limiting Assumption 2.1. Consider a state $s\in\mathcal{S}$ and three arms $a_{i},a_{j},a_{\tilde{j}}$ such that
@@
By our selection of $a_{j},a_{\tilde{j}}$, we know that $\boldsymbol{p}_{i,j}(i)-\boldsymbol{p}_{i,{\tilde{j}}}(i)\leq 0$; hence, the third term is non-positive. Interestingly, as we show in Claim 5 in Section H, stochastic order does not imply that $W^{\star}(s\setminus\{a_{j}\})\geq W^{\star}(s\setminus\{a_{\tilde{j}}\})$; thus, it is unclear that the expression
@@
which appears in Inequality (5), is non-negative. Therefore, we cannot prove Inequality (5) without revealing the structure of $W^{\star}$, even when Assumption 2.1 holds.
@@ 4.1 The Case of $T\rightarrow\infty$
An immediate corollary of Observations 4.1 and 4.2 is a sandwich argument. For an optimal policy $\pi^{\star}$, namely, $W(\pi^{\star},A)=W^{\star}(A)$, it holds that
@@
Corollary 4.1.
Fix any arbitrary IRSR instance satisfying Assumption 2.1. It holds that
@@
Input: the OGP policy
1 $s\leftarrow A$
2 while ${\small\textnormal{{OGP}}}(s)\neq\emptyset$ do
      // $s$ is not a terminal state
3      select ${\small\textnormal{{OGP}}}(s)$, and denote the realized action by $a_{k}$
4       $s\leftarrow s\setminus\{a_{k}\}$
5if $x_{a_{k}}>0$ for some $a_{k}\in A$ then
      // A positive reward is revealed
6      $a_{k^{\star}}\leftarrow\operatorname*{arg\,max}_{a_{i}}x(a_{i})$
       // best among all the explored arms
7      execute Bernoulli trials mixing $a_{k^{\star}}$ with every other unexplored arm until all are revealed
8       select the best arm forever
9
10else
11       select the default arm $a_{0}$ forever
12
ALGORITHM 2 Individually Rational Exploration via GMDP and Bernoulli Trials (IREGB)
@@ 4.2 Convergence Rate
We begin by noting that in special cases, IREGB can be slightly modified to achieve ${\small\textnormal{{OPT}}}_{\infty}$ even for a finite $T$. To illustrate, consider $(X(a_{i}))_{i}$ that are supported on $\{x^{-},x^{+}\}$, with $x^{+}>0$ and $x^{-}<0$. In such a case, revealing a positive reward of $x^{+}$ suggests that we need not explore any further (since the other rewards cannot outperform $+1$); thus, the Bernoulli trials in Line 2 become redundant. Formalizing this intuition,
Proposition 4.1.
@@
Fix any arbitrary IRSR instance such that $(X(a_{i}))_{i}\in\{x^{-},x^{+}\}$ w.p. 1, i.e., the rewards take only two possible values almost surely. Let ${\small\textnormal{{IREGB}}}^{\prime}$ be a modified version of IREGB that exploits once a reward of $x^{+}$ is realized. Then, there exists $T_{0}$ such that whenever $T\geq T_{0}$, $SW_{T}({\small\textnormal{{IREGB}}}^{\prime})={\small\textnormal{{OPT}}}_{\infty}$.
@@
However, IREGB cannot achieve ${\small\textnormal{{OPT}}}_{\infty}$ in finite time due to two factors. First, the Bernoulli trials in Line 2 might take several rounds, thereby delaying the potentially higher rewards hidden by risky arms. Consequently, the number of rounds that can benefit from these high rewards is limited (unlike in the $T\rightarrow\infty$ case). Second, the Bernoulli trials can be unfruitful in expectation. To illustrate, assume that there are only a few rounds left and one unexplored arm from $\textnormal{neg}(A)$. In such a case, the one-time cost of exploring this arm outweighs the value it would provide in the remaining rounds. Deciding whether to mix such arms in Bernoulli trials or to exploit the best-seen arm constitutes another barrier. We leave this technical challenge for future work. Instead, our goal is to bound the time it takes to reveal all arms and address the potential performance gap it causes.
@@
Given an IRSR instance, let the random variable $\delta$ denote the highest positive reward among the arms of $\textnormal{pos}(A)$, i.e., $\delta=\max_{a_{i}\in\textnormal{pos}(A):X(a_{i})>0}X(a_{i})$. Recall that in Line 2 of Algorithm 2, we execute Bernoulli trials with the best-seen arm. The term $\delta$ constitutes a lower bound on the reward of that arm $a_{k^{\star}}$. Furthermore, let $\eta=\max_{a_{i}\in A:\mu(a_{i})<0}\left|\mu(a_{i})\right|$. Notice that $\eta$ quantifies the highest absolute value of an arm in $\textnormal{neg}(A)$. The success probability of each Bernoulli trial, which leads to exploring one additional arm, is at least $\frac{\delta}{\delta+\eta}$. Consequently, after $K(1+{\eta}\operatorname{\mathbb{E}}\left[{\frac{1}{\delta}}\right])$ Bernoulli trials in expectation, we would explore all arms. We formalize this intuition through Proposition 4.2 below.
Proposition 4.2.
@@
Fix any arbitrary IRSR instance satisfying Assumption 2.1, and let $\delta>0$ and $\eta$ be the quantities defined above. There exists $T_{0}$ such that whenever $T>T_{0}$, it holds that $SW_{T}({\small\textnormal{{IREGB}}})\geq\left(1-\frac{K(1+\eta\cdot\operatorname{\mathbb{E}}\left[{\frac{1}{\delta}}\right]}{T}\right){\small\textnormal{{OPT}}}_{\infty}.$
@@
To illustrate, assume that $(X(a_{i}))_{i}$ are arbitrarily distributed in the discrete set $\{-H,\dots,H\}$. In such a case, as long as $\delta>0$, $\frac{\eta}{\delta}\leq H$ holds almost surely; thus, Proposition 4.2 suggests that IREGB is optimal up to a multiplicative factor of $\frac{K(H+1)}{T}$ or an additive factor of $\frac{KH(H+1)}{T}$.

### v2

@@ 2.2 Stochastic Order
Assumption 1
The rewards of the arms in $\textnormal{neg}(A)$ are stochastically ordered.
Notice that this assumption applies only to a priori inferior arms, i.e., the arms in $\textnormal{neg}(A)$, and not to all arms. There are many natural cases for this assumption: Unit-variance (or any fixed variance) Gaussian, as in Example 1, Bernoulli, log-normal, truncated normal, etc. When a formal statement relies on Assumption 1, we mention it explicitly.
@@ 3 An Auxiliary Goal Markov Decision Process
3.1 Motivation for GMDP: Bernoulli Trials
@@
We call mixing an observed arm with a positive realized reward and an unobserved arm with a negative expected reward a Bernoulli Trial. Observation 1 suggests that if we discover an arm with a positive reward, we can use Bernoulli trials to reveal all arms in finite time almost surely.
To illustrate, recall Example 1. Assume that we are in the third round, have already explored $a_{1}$ and $a_{2}$, and discovered that $X(a_{1})=x(a_{1})>0$ and that $X(a_{2})<0$. Furthermore, we have not yet explored $a_{3}$ and $a_{4}$ (one of which may still hide the highest realized reward among all arms). Consider the portfolio
@@
thus, $\boldsymbol{p}$ is MIR. If we pick this portfolio, Nature flips coins to choose either $a_{1}$ or $a_{3}$, with the latter being selected with positive probability. This Bernoulli trial might result in selecting $a_{1}$, but we can repeat it until Nature picks $a_{3}$. The number of rounds required to explore $a_{3}$ follows the Geometric distribution with a success probability of $\nicefrac{{x(a_{1})}}{{x(a_{1})-\mu(a_{3})}}$ in each Bernoulli trial. Indeed, by executing Bernoulli trials until the first success, we can guarantee exploring $a_{3}$. Zooming out from Example 1 to the general case, Observation 1 suggests that once a positive reward is realized, we can execute enough Bernoulli trials to explore all the arms in finite time.
Observation 1 calls for modeling that abstracts the setting once a positive reward is realized. More specifically, we can focus on the case where all unexplored arms are revealed instantly after a positive reward is discovered. To that end, we analyze the induced constrained Goal Markov Decision Process (GMDP), which we present in the next subsection. Our goal is to find the optimal policy for this GMDP, and then translate it to an asymptotically optimal algorithm for the corresponding IRSR instance. [^4] We keep the terms mechanism and algorithm for solutions to IRSR and use the term policy for solutions of the GMDP.
3.2 Constructing the GMDP
@@
  • In every state $s$, the feasible actions are the set of MIR portfolios w.r.t. the prior information, i.e., from $\textit{MIR}(s)=\left\{\boldsymbol{p}\in\Delta(s):\sum_{a\in s}\boldsymbol{p}(a)\mu(a)\geq 0\right\},$ where $\Delta(s)$ is the set of all distributions over the elements of $s$. If $\textit{MIR}(s)$ is empty, then we say that $s$ is terminal.
@@
    $R(s)\stackrel{{\scriptstyle\text{def}}}{{=}}\begin{cases}\max_{a\in A}X(a)&\textnormal{if $\textit{MIR}(s)=\emptyset$ and $\max_{a^{\prime}\in A\setminus s}X(a^{\prime})>0$}\\ 0&\textnormal{otherwise}\end{cases}.$
@@
A (stationary) policy is a mapping from states to portfolios. Due to standard arguments, which are further elaborated in Section 0.A, there exists an optimal policy that is also stationary. Hence, we limit our attention to stationary policies only. Let $W(\pi,s)$ be the state-value function: The expected reward of a policy $\pi$ starting at the initial state $s$. Namely,
$W(\pi,s)=\begin{cases}R(s)&\textnormal{if }\textit{MIR}(s)=\emptyset\\ \sum_{a\in s}\Pr(s\setminus\{a\}\mid s,\pi(s))W(\pi,s\setminus\{a\})&\textnormal{otherwise}\end{cases}.$ (3)
@@ 3.3 $\mathcal{P}$-valid Policies and an Inefficient Solution via Dynamic Programming
mixes $a_{i}$ and $a_{j}$ while maximizing the probability of exploring $a_{j}$ (the a priori inferior arm). The reader can verify that $\boldsymbol{p}_{i,j}$ is indeed MIR, yielding an expected value of precisely zero. For completeness, for every $a_{i}\in\textnormal{pos}(A)$, we also define $\boldsymbol{p}_{i,i}$ as a deterministic selection of $a_{i}$, e.g., $\boldsymbol{p}_{i,i}(a)=1$ if $a=a_{i}$, and zero otherwise. Next, we define $\mathcal{P},\mathcal{P}^{\prime}$ such that
@@
Definition 3 ($\mathcal{P}$-valid Portfolio)
A portfolio $\boldsymbol{p}$ is $\mathcal{P}$-valid with respect to a state $s\in\mathcal{S}$ if
  • $\boldsymbol{p}\in\textit{MIR}(s)$, and
@@
We say a policy is $\mathcal{P}$-valid if, for all states $s\in\mathcal{S}$, $\pi(s)$ is a $\mathcal{P}$-valid portfolio. Notice that $\mathcal{P}$-valid policies are MIR by definition. Furthermore, they mix at most two actions from $A$ in each state; hence, there are $O(K^{2})$ actions $\mathcal{P}$-valid policies can take in every state. This is in sharp contrast to the set of actions $\textit{MIR}(s)$ for a state $s$, which is generally a convex polytope with infinitely many actions.
Due to the linearity of the value function $W(\pi,s)$ in $\pi(s)$, the GMDP exhibits a nice structural property, as captured by the following Proposition 1.
Proposition 1
@@
The proof of this proposition appears in Section 0.H. While the GMDP is defined using a continuous action space in each state (the convex polytope $\textit{MIR}(s)$), Proposition 1 allows us to narrow down the search to $\mathcal{P}$-valid portfolios. Particularly, we can find an optimal policy with dynamic programming using a bottom-up approach in time $O(2^{K}K^{2})$, as we explain next.
@@ 3.4 Efficiently Computing an Optimal GMDP Policy
0:  a state $s\subseteq A$
0:  a $\mathcal{P}$-valid portfolio or $\emptyset$
1:  if $s$ is terminal  then
2:     return $\emptyset$
3:  else
4:     pick any arbitrary $a_{i}\in\textnormal{pos}(s)$
5:     if $\textnormal{neg}(s)=\emptyset$ then
6:        return  $\boldsymbol{p}_{i,i}$
7:     else
8:        pick $a_{j^{\star}}\in\operatorname*{arg\,max}_{a_{j}\in\textnormal{neg}(s)}\mu(a_{j})$
9:        return  $\boldsymbol{p}_{i,j^{\star}}$
We start this subsection by presenting an optimal GMDP policy (hereinafter OGP for shorthand), which we formalize via Policy 1, and argue for its optimality later. Given a state $s$, ${\small\textnormal{{OGP}}}(s)$ operates as follows. If $s$ is terminal, it returns the empty set (Lines 12). Otherwise, if $s$ is non-terminal, we enter the “else” clause in Line 3, and pick any arbitrary arm $a_{i}$ from $\textnormal{pos}(s)$. Then, we have two cases. If $\textnormal{neg}(s)$ is empty, OGP returns $\boldsymbol{p}_{i,i}$ (Line 6). Else, it picks the best arm from $\textnormal{neg}(s)$ in terms of expected reward, which we denote by $a_{j^{\star}}$ (Line 8), and returns $\boldsymbol{p}_{i,j^{\star}}$ (Line 9). Overall, OGP returns $\emptyset$ if $s$ is terminal or a $\mathcal{P}$-valid portfolio if $s$ is non-terminal.
Theorem 3.1
Fix any arbitrary IRSR instance satisfying Assumption 1 and observe its induced GMDP instance. For any state $s\subseteq A$, OGP prescribed in Policy 1 satisfies $W({\small\textnormal{{OGP}}},A)=W^{\star}(A)$.
Proof (Proof Overview of Theorem 3.1)
We defer the full proof to Section 0.A and outline it below. To prove this theorem, we focus on the dynamic programming explained at the end of the previous subsection. We unveil its crux and flesh out the index structure OGP uses. A crucial ingredient of the analysis is the probability of reaching the empty state, representing the case in which we explored all arms. More formally, for every state $s\in\mathcal{S}$ and a policy $\pi$, we denote by $Q(\pi,s)$ the probability starting at $s\subseteq A$, following the policy $\pi$ and reaching the empty state. Note that, similarly to $W$ the state-value function, $Q$ is defined recursively: Namely, if $\pi(s)=\boldsymbol{p}_{i,j}$ for a non-terminal state $s$, then
@@
Lemma 1 (Equivalence Lemma)
@@
The Equivalence Lemma asserts that $Q$ is policy-independent, a property that significantly simplifies the analysis. The proof of this lemma is intricate and relies on a careful factorization of the elements of $Q$. Interestingly, the arguments we have mentioned so far, including the Equivalence Lemma, are independent of Assumption 1.
Equipped with the Equivalence Lemma, we use the recursive structure of $Q$ and the reward function $W$ to prove the optimality of OGP inductively over the sizes of $\textnormal{pos}(s)$ and $\textnormal{neg}(s)$. This step makes use of further insights into the dynamic programming procedure, such as monotonicity in the set of available arms.
@@
Figure 1: Illustration of OGP for Example 1.
@@ 3.5 Stochastic Order and Non-triviality
Before leveraging OGP to an optimal IRSR algorithm, it is important to describe why Theorem 3.1 is non-trivial to prove, even with the limiting Assumption 1. Consider a state $s\in\mathcal{S}$ and three arms $a_{i},a_{j},a_{\tilde{j}}$ such that
@@
By our selection of $a_{j},a_{\tilde{j}}$, we know that $\boldsymbol{p}_{i,j}(i)-\boldsymbol{p}_{i,{\tilde{j}}}(i)\leq 0$; hence, the third term is non-positive. Interestingly, as we show in Claim 5 in Section 0.H, stochastic order does not imply that $W^{\star}(s\setminus\{a_{j}\})\geq W^{\star}(s\setminus\{a_{\tilde{j}}\})$; thus, it is unclear that the expression
@@
which appears in Inequality (5), is non-negative. Therefore, we cannot prove Inequality (5) without revealing the structure of $W^{\star}$, even when Assumption 1 holds.
@@ 4.1 The Case of $T\rightarrow\infty$
An immediate corollary of Observations 2 and 3 is a sandwich argument. For an optimal policy $\pi^{\star}$, namely, $W(\pi^{\star},A)=W^{\star}(A)$, it holds that
@@
Corollary 1
Fix any arbitrary IRSR instance satisfying Assumption 1. It holds that
@@
Algorithm 1 MIR Exploration via GMDP and Bernoulli Trials (IREGB)
0:  the OGP policy
1:  $s\leftarrow A$
2:  while ${\small\textnormal{{OGP}}}(s)\neq\emptyset$ do #$s$ is not a terminal state
3:     select ${\small\textnormal{{OGP}}}(s)$, and denote the realized action by $a_{k}$
4:     $s\leftarrow s\setminus\{a_{k}\}$
5:  if $x_{a_{k}}>0$ for some $a_{k}\in A$ then #A positive reward is revealed
6:     $a_{k^{\star}}\leftarrow\operatorname*{arg\,max}_{a_{i}}x(a_{i})$best among all the explored arms
7:     execute Bernoulli trials mixing $a_{k^{\star}}$ with every other unexplored arm until all are revealed
8:     select the best arm forever
9:  else, select the default arm $a_{0}$ forever
@@ 4.2 Convergence Rate
We begin by noting that in special cases, IREGB can be slightly modified to achieve ${\small\textnormal{{OPT}}}_{\infty}$ even for a finite $T$. To illustrate, consider $(X(a_{i}))_{i}$ that are supported on $\{x^{-},x^{+}\}$, with $x^{+}>0$ and $x^{-}<0$. In such a case, revealing a positive reward of $x^{+}$ suggests that we need not explore any further (since the other rewards cannot outperform $+1$); thus, the Bernoulli trials in Line 7 become redundant. Formalizing this intuition,
Proposition 2
@@
Fix any arbitrary IRSR instance such that $(X(a_{i}))_{i}\in\{x^{-},x^{+}\}$ w.p. 1, i.e., the rewards are almost surely supported on only two values. Let ${\small\textnormal{{IREGB}}}^{\prime}$ be a modified version of IREGB that exploits once a reward of $x^{+}$ is realized. Then, there exists $T_{0}$ such that whenever $T\geq T_{0}$, $SW_{T}({\small\textnormal{{IREGB}}}^{\prime})={\small\textnormal{{OPT}}}_{T}$.
@@
However, IREGB cannot achieve ${\small\textnormal{{OPT}}}_{\infty}$ in finite time due to two factors. First, the Bernoulli trials in Line 7 might take several rounds, thereby delaying the potentially higher rewards hidden by risky arms. Consequently, the number of rounds that can benefit from these high rewards is limited (unlike in the $T\rightarrow\infty$ case). Second, the Bernoulli trials can be unfruitful in expectation. To illustrate, assume that there are only a few rounds left and one unexplored arm from $\textnormal{neg}(A)$. In such a case, the one-time cost of exploring this arm outweighs the value it would provide in the remaining rounds. Deciding whether to mix such arms in Bernoulli trials or to exploit the best-seen arm constitutes another barrier. We leave this technical challenge for future work. Instead, our goal is to bound the time it takes to reveal all arms and address the potential performance gap it causes.
@@
Given an IRSR instance, let the random variable $\delta$ denote the highest positive reward among the arms of $\textnormal{pos}(A)$, i.e., $\delta=\max_{a_{i}\in\textnormal{pos}(A):X(a_{i})>0}X(a_{i})$. Recall that in Line 7 of Algorithm 1, we execute Bernoulli trials with the best-seen arm. The term $\delta$ constitutes a lower bound on the reward of that arm $a_{k^{\star}}$. Furthermore, let $\eta=\max_{a_{i}\in A:\mu(a_{i})<0}\left|\mu(a_{i})\right|$. Notice that $\eta$ quantifies the highest absolute value of an arm in $\textnormal{neg}(A)$. The success probability of each Bernoulli trial, which leads to exploring one additional arm, is at least $\frac{\delta}{\delta+\eta}$. Consequently, after $K(1+{\eta}\operatorname{\mathbb{E}}\left[{\frac{1}{\delta}}\right])$ Bernoulli trials in expectation, we would explore all arms. We formalize this intuition through Proposition 3 below.
Proposition 3
@@
Fix any arbitrary IRSR instance satisfying Assumption 1, and let $\delta>0$ and $\eta$ be the quantities defined above. There exists $T_{0}$ such that whenever $T>T_{0}$, it holds that $SW_{T}({\small\textnormal{{IREGB}}})\geq\left(1-\frac{K(1+\eta\cdot\operatorname{\mathbb{E}}\left[{\frac{1}{\delta}}\right]}{T}\right){\small\textnormal{{OPT}}}_{\infty}.$
@@
To illustrate, assume that $(X(a_{i}))_{i}$ are arbitrarily distributed in the discrete set $\{-H,\dots,H\}$. In such a case, as long as $\delta>0$, $\frac{\eta}{\delta}\leq H$ holds almost surely; thus, Proposition 3 suggests that IREGB is optimal up to a multiplicative factor of $\frac{K(H+1)}{T}$ or an additive factor of $\frac{KH(H+1)}{T}$.